OpenStudy (anonymous):

What is the domain of f(x)=2x-1/ x^2+2x-3

OpenStudy (anonymous):

it is \[\frac{2x-1}{x^2+3x-3}\]?

OpenStudy (anonymous):

no the bottom is x^2+2x-3

OpenStudy (anonymous):

if so, ignore the numerator and find the zeros of the denominator. \[x^2+2x-3=0\] \[(x-1)(x+3)=0\] \[x=1\] or\[x=-3\] so all real numbers except 1 and -3